Cn you please help me solve this system of equations?
3x-y=5
2x-3y=-6
: Cn you please help me solve this system of equations?
3x-y=5
2x-3y=-6

Answer by stanbon(19008)

3x-y=5
2x-3y=-6
--------------
Multiply thru the 1st equation by 3 to get:
9x - 3y = 15
2x - 3y = -6
----------------
Subtract the 2nd equation from the 1st to get:
7x = 21
x = 3
-----------------
Substitute into 3x-y = 5 to solve for Y:
3*3 -y = 5
y = 4
------------
Final solution:
x =3; y = 4
================
Cheers,
Stan H.